Assume our velocity selector is a bit leaky and lets the other 2 isotopes through to the mass spectrometer. 3. What is the spacing between the isotopes when they strike the detectors?

Just 3 and 4An Iron-Nickel sample from a meteorite was placed into a velocity selector and mass spectrometer. Experimenters expect 3 prim
Answer #1

Part-(1)

(A) Direction of electric field in velocity selector is considered as direction of +z-axis as shown in figure (orange arrow lines )

(B) Positive ions deflected upwards in the assumed electric field .

If velocity of +ve ions is in the +y axis direction , in order to get downward deflection to balance the deflection due to electric field, magnetic field has to be applied in +x axis direction as shown in figure.

Magnetic field direction is out of page,shown as \odot in figure

(C) Circular path of +ve ions in mass spectrometer region is shown as blue line

---------------------------------------------------

Part-(2)

Magnetic field induction B applied to velocity selector is calculated from following equation

E = v B

where E is electric field intensity and v is speed of ions

Since ions are accelerated by 3300 V , speed of ions is calculated as follows

eV = (1/2) m v2

where e is charge on singly ionised 58Ni , V is accelerating potential and m is amss of ion

v = \sqrt{\frac{2 e V}{m}} = \sqrt{\frac{2 \times 1.602 \times 10^{-19}\times 3300}{58\times 1.66\times 10^{-27}}} \approx 10^{5} m/s

Hence magentic field induction B = E / v = 3300 / ( 1.1 \times 10-2\times 105 ) = 3 T

--------------------------------------------------

Part-(3)

Radius of singly ionised 58Ni ion is claculated by equating magnetic force and centripetal force as follows

q v B = (m v2 ) / R   

where R is radius of circular path .

We get radius R = ( m v ) / ( q B )   ..................(1)

R = ( 58 \times 1.66 \times 10-27\times 105 ) / ( 1.602 \times 10-19\times 3 ) = 20 \times 10-3 m

Radius R = 20 mm

We see from eqn.(1), Radius directly proportional to mass ,

Radius of 56Fe= 20 \times (56/58) = 19.34 mm

Radius of 60Ni= 20 \times (60/58) = 20.69 mm

Hence spacing between 58Ni and 56Fe = 2 \times ( 20 - 19.34) mm = 1.32 mm

Spacing between 58Ni and 60Ni = 2 \times ( 20.69 - 20) mm = 1.38 mm

----------------------------------------------------------------------------------------------------

If we see from eqn.(1), Radius inversely proportional to charge q and directly proportional to mass.

Hence for doubly ionised ion, radius will be halved.

So we may not be knowing if the detected ion is doubly ionised or having half of mass.